A variable line makes intercepts on the coordinate axes the sum of whose squares is constant and is equal to a^2. Find the locus of the foot of perpendicular from the origin to this line.

Grade:11

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Karthik,

 

eq of line is x/p + y/q = 1 where p , q are intercept on x. y axis

alos given,  p2 + q2 = a2      eq 1

let foot of perpendicular is x1, y1

 

this point should lie on x/p + y/q = 1

so,  x1/p + y1 /q =1

 

also  (y1/x1 ) * ( - q/p ) = -1       eq2

find value of p, q from eq1 and eq2 and sunstitute in x1/p + y1 /q =1.

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay